OpenStudy (anonymous):

simplify (3z^3)^4 write without parentheses

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

same as before you distribute the exponent, thus getting rid of the parentheses :)

OpenStudy (darkbluechocobo):

3^4 z^12

OpenStudy (darkbluechocobo):

You distribute 4 to each

OpenStudy (darkbluechocobo):

i mean ^4*

OpenStudy (jdoe0001):

hmmm I see...well still the same more or less

OpenStudy (anonymous):

so 3 just stays 3?

OpenStudy (jdoe0001):

\(\large \bf (a^n)^m\implies a^{n\cdot m} \\ \quad \\ \quad \\ (3z^3)^4\implies 3^4(z^3)^4\implies 3^4z^{3\cdot 4}\implies 3^4z^{12}\)

OpenStudy (darkbluechocobo):

No 3^4 so 3x3x3x3

OpenStudy (anonymous):

ok what about simplify (-3a^3/b)^3 write answer without parentheses

OpenStudy (darkbluechocobo):

With exponents you add ^3 to the current exponent

OpenStudy (jdoe0001):

\(\bf \left(-\cfrac{3a^3}{b}\right)^3 ?\)

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (darkbluechocobo):

multiplying

OpenStudy (darkbluechocobo):

* so a^3*3

OpenStudy (darkbluechocobo):

(3x3x3)(a^3*3)/b^3

OpenStudy (anonymous):

I don't understand :(

OpenStudy (darkbluechocobo):

Ok so have you dealt with distribution property before?

OpenStudy (anonymous):

yes

OpenStudy (darkbluechocobo):